# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v1

@@ 1 Introduction
  • We identify a misalignment between standard fine-tuning with CE loss and the pass@N coverage metric at test time. (Section 4.1)
  • We develop and experimentally verify a framework that suggests this misalignment arises from overconfidence induced by training on CE loss. (Sections 4.2 and 4.3)
  • We propose an alternate loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers in MATH and MiniF2F. (Sections 5.1, 5.2 and 5.3)
  • We extend our algorithm to more complex test-time scenarios including search over proof-trees of varying shapes to improve automated theorem proving, and answering math questions using Chain-of-Thought reasoning traces, demonstrating improved mathematical reasoning in both cases. (Sections 5.3 and 5.4)
@@ 2 Related Works
Test-time compute. Multiple strategies have been proposed for improving LLM performance by scaling test-time compute. Jones (2021) demonstrates a tradeoff between train- and test-time compute in the toy model of a board game. OpenAI’s O1 model demonstrates remarkable performance gains when scaling test-time compute (OpenAI, 2024). Closely related to our work, Brown et al. (2024) observed an exponentiated power law between coverage and the number of samples for in-context-learning evaluation. Snell et al. (2024), in turn, have explored compute-optimal strategies for effectively scaling test-time compute.
@@
Post-training for mathematical reasoning. Similarly, multiple post-training techniques have been proposed to improve mathematical reasoning in LLMs. Instruction-tuning and reinforcement learning with human feedback have been shown to boost model performance on math (Yue et al., 2024; Lightman et al., 2024; Uesato et al., 2022), while continued training on math- or code-specific domain data enhances models’ reasoning abilities for downstream mathematical tasks (Lewkowycz et al., 2022; Azerbayev et al., 2024; Yang et al., 2024; Shao et al., 2024; Ying et al., 2024). Rejection-sampling (Zelikman et al., 2022) and self-improvement techniques (Qi et al., 2024), in turn, are useful to augment the training data for SFT. More recent approaches (OpenAI, 2024; DeepSeek-AI: Daya Guo et al., 2025) have incorporated reinforcement learning and achieved exceptional reasoning capabilities in various domains, such as math and coding. Although our paper primarily focuses on supervised fine-tuning to enhance pretrained models’ math capabilities, our loss function can be applied to other settings that train under CE loss, such as continual training, instruction-tuning, and data augmentation.
@@
Data pruning and hard example mining. The interpretation of the loss function we derive below can be related to data pruning and hard example mining. Data selection is often applied to curate high-quality datasets for pretraining (Marion et al., 2023), where Sorscher et al. (2022) shows that pruning easy samples can improve pretraining loss scaling as a function of dataset size. On the other hand, hard example mining focuses on identifying and emphasizing challenging samples to improve model performance (Shrivastava et al., 2016). In the domain of mathematical reasoning, Tong et al. (2024) found a difficulty imbalance in rejection-sampled datasets and showed that more extensive training on difficult samples improves model performance.
@@ 4.1 The CE loss induces overfitting for pass@N
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than LLama-3-8B-Instruct to avoid potential leakage of the MATH dataset into LLama-3-8B-Instruct through post-training. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer, without a reasoning trace. We will discuss training with CoT in Section 5.4 for MATH reasoning traces.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
To demonstrate more generally the existence of tradeoffs between confident exploitation of a few answers versus unconfident exploration of many answers, as a function of the number of passes $N$, we prove two lemmas. To set up these lemmas, given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Thus $\hat{p}_{1}$ is the model’s maximal confidence across all answers. Moreover, let $p_{i}$ be the probability across all the problems that the $i^{\text{th}}$ ranked answer is actually correct. Assume the model policy is approximately well calibrated so that higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$. Indeed, optimal policies maximizing pass@N coverage in Equation 2 are approximately well calibrated (See Lemma A.1). $p_{1}$ is then the model’s maximal accuracy across all answers. We prove:
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas 4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure 1 (leftmost). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure 1 (left) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance.
@@ 5.1 A solution that naturally prevents overconfidence
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure 2). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure 1 (leftmost).
@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
Figure 3: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
@@ 5.4 Approximate DCO also improves MATH answering with chain-of-thought reasoning
Figure 5: Overconfidence persists in CoT fine-tuning with CE loss, which DCOa successfully limits. We fine-tune a Llama-3-8B base model on the MATH dataset on CoT traces. We plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. Left: for CE loss, the model becomes more confident in its most likely answers, as the confidence distribution shifts slightly to the right as training progresses. This effect is milder than in the direct answer setting (Figure 1). Right: for DCOa loss, the rightward shift in the confidence distribution of the most likely answer over training time is limited; DCOa again limits overconfidence relative to CE loss.
@@ Acknowledgments
We would like to thank Zhinan Cheng, Clémentine Dominé, Marco Fumero, David Klindt, Daniel Kunin, Ben Sorscher, Atsushi Yamamura for helpful discussions. S.G. thanks the James S. McDonnell and Simons Foundations, NTT Research, Schmidt Foundation and an NSF CAREER Award for support. S.D. and F.C. were partially supported by the McKnight Foundation, the Simons Foundation and an NIH CRCNS grant R01DC020874.
@@ Appendix B Experimental details
All experiments are performed on machines with 8 NVIDIA H100 GPUs or 8 NVIDIA A100 GPUs. Our codebase uses PyTorch (Ansel et al., 2024), Accelerate (Gugger et al., 2022), and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm (Kwon et al., 2023) for efficient inference. Code will be made available here.
MATH. For experiments with MATH dataset, we fine-tune the Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. In Sections 4 and 5.2 and Figures 1 and 3, we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section 5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
@@
Appendix C Addtional results
C.1 Theorem proving

### v2

@@ 1 Introduction
  • We identify a misalignment between standard fine-tuning with CE loss and the pass@N coverage metric at test time (Section 4.1).
  • We develop and experimentally verify a framework that suggests this misalignment arises from overconfidence induced by training on CE loss (Sections 4.2 and 4.3).
  • We propose an alternate loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers in MATH and MiniF2F (Sections 5.1, 5.2 and 5.3).
  • We extend our algorithm to more complex test-time scenarios including searching over proof-trees of varying shapes to improve automated theorem proving, and answering math questions using Chain-of-Thought reasoning traces, demonstrating improved mathematical reasoning in both cases (Sections 5.3 and 5.4).
@@ 2 Related Works
Test-time compute and pass@N strategy. A growing body of work explores how training and test-time strategies interact to affect model performance. Jones (2021) highlights a tradeoff between train and test-time compute in a controlled board game setting. Closely related to our work, Brown et al. (2024) identify an exponentiated power law between coverage and the number of in-context samples. Snell et al. (2024), in turn, explore compute-optimal strategies for scaling test-time compute. Our paper focuses primarily on the pass@N test-time strategy. Gui et al. (2024) show that, in the best-of-N setting, the sample distribution is nearly optimally aligned with the reward model. To mitigate the excessive test-time compute costs associated with best-of-N sampling, Gui et al. (2024); Sessa et al. (2024) propose alignment algorithms to distill the sample distribution. Li et al. (2024) argue that fine-tuning with cross-entropy loss can limit output diversity and propose a maximum entropy-based method to address this issue. Our work similarly identifies limitations of cross-entropy when scaling test-time compute, but focuses on the alignment of training and test-time strategies. Zhang et al. (2024) explore methods for preventing catastrophic forgetting in supervised fine-tuning over diverse domains.
@@
Post-training for mathematical reasoning. Various post-training techniques have been proposed to improve mathematical reasoning in LLMs. Instruction-tuning and reinforcement learning with human feedback have been shown to boost model performance on math (Yue et al., 2024; Lightman et al., 2024; Uesato et al., 2022), while continued training on math- or code-specific domain data enhances models’ reasoning abilities for downstream mathematical tasks (Lewkowycz et al., 2022; Azerbayev et al., 2024; Yang et al., 2024; Shao et al., 2024; Ying et al., 2024). Rejection-sampling (Zelikman et al., 2022) and self-improvement techniques (Qi et al., 2024), in turn, are useful to augment the training data for SFT. More recent approaches (OpenAI, 2024; DeepSeek-AI: Daya Guo et al., 2025) have incorporated reinforcement learning and achieved exceptional reasoning capabilities in various domains, such as math and coding. Although our paper primarily focuses on supervised fine-tuning to enhance pretrained models’ math capabilities, our loss function can be applied to other settings that train under CE loss, such as continual training, instruction-tuning, and data augmentation.
@@
Data pruning and hard example mining. The interpretation of the loss function we derive below can be related to data pruning and hard example mining (see discussion of sample difficulty in Section 4.4). Data selection is often applied to curate high-quality datasets for pretraining (Marion et al., 2023), where Sorscher et al. (2022) shows that pruning easy samples can improve pretraining loss scaling as a function of dataset size. Zhou et al. (2023); Ye et al. (2025) demonstrate that supervised fine-tuning on even a very small dataset can yield remarkably strong performance in alignment and reasoning tasks. On the other hand, hard example mining focuses on identifying and emphasizing challenging samples to improve model performance (Shrivastava et al., 2016). In the domain of mathematical reasoning, Tong et al. (2024) find a difficulty imbalance in rejection-sampled datasets and show that more extensive training on difficult samples improves model performance.
@@ 4.1 The CE loss induces overfitting for pass@N
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than LLama-3-8B-Instruct to avoid potential leakage of the MATH dataset into LLama-3-8B-Instruct through post-training. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer, without a reasoning trace. We will discuss training with CoT in Section 5.4 for MATH reasoning traces. We confirm that the results presented in this section also hold for a larger model, Llama-3-70B-base, and refer readers to Section C.2 for additional results.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
To demonstrate more generally the existence of tradeoffs between confident exploitation of a few answers versus unconfident exploration of many answers, as a function of the number of passes $N$, we prove two lemmas. To set up these lemmas, given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Thus $\hat{p}_{1}$ is the model’s maximal confidence across all answers. Moreover, let $p_{i}$ be the probability across all the problems that the $i^{\text{th}}$ ranked answer is actually correct. Assume the model policy is approximately well calibrated so that higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$. We empirically verify that the model trained with CE loss on the MATH dataset approximately satisfies this assumption (Figure 14). Indeed, optimal policies maximizing pass@N coverage in Equation 2 are approximately well calibrated (See Lemma A.1). $p_{1}$ is then the model’s maximal accuracy across all answers. We prove:
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas 4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure 1 (leftmost). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure 1 (left) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance. In Section C.3, we observe a similar trend of overconfidence on the out-of-distribution test set AIME24, further highlighting the generality of this failure mode.
4.4 Easy data drives overconfidence
Figure 2: Easy data drives overconfidence and degrades performance when scaling test-time compute. For the experiments in Figure 1, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ at two points during the first training epoch: 11% (early) and 86% (late). At each stage, the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$ is evaluated on batches of previously unseen training data of each difficulty level defined in the MATH dataset. Early in training, data from all difficulty levels contribute to decreasing the test loss (left). However, later in training, easier examples (difficulty level 2) provide no further benefit, while the easiest examples (difficulty level 1) actively degrade test performance (right). The plotted $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is an average over batches of unseen data, and we use $N=256$, corresponding to pass@256.
We have shown above that overconfidence limits performance gains from scaling test-time compute. Here we investigate whether all training data contribute equally to the drop in pass@N test performance. To quantify this effect, we define the test loss as the negative log-probability of coverage, $\mathcal{L}_{\text{test}}^{N}=-\sum_{(x,y)\in\mathcal{D}^{\mathrm{test}}}\mathcal{\log}\ \mathcal{C}_{(x,y)}^{N}$, where $\mathcal{C}_{(x,y)}^{N}$ denotes the coverage for a single test example $(x,y)$. To measure how a training batch $\mathcal{B}$ influences the final test loss, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ for the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$, where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example. A negative value, $\nabla_{g}\mathcal{L}_{\text{test}}^{N}<0$, indicates that a gradient step in the direction of $g$ decreases the test loss.
To analyze the impact of data difficulty, we group training data by difficulty level (as specified in the MATH dataset) and use the directional derivative to examine how a batch of previously unseen data from a given difficulty level would affect $\mathcal{L}_{\text{test}}^{N}$ at different stages of training. Early in training, we observe that $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is negative across all difficulty levels, indicating that data from all difficulty levels contribute positively to reducing the test loss (Figure 2, left). However, near the end of the first epoch, training on easier examples (difficulty levels 1 and 2) no longer improves performance, and the easiest examples (level 1) actively degrade pass@N test performance (Figure 2, right). This shows how continued training on easy data can harm test performance when scaling test-time compute.
@@ 5.1 A solution that naturally prevents overconfidence
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure 3). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure 1 (leftmost). Indeed, $F(N,\hat{p}(y|x))$ is smaller for easier examples (Figure 15, right), effectively regularizing their contribution. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section 4.4.
@@ 5.2 DCO can prevent overconfidence and rescue test-time scaling
Figure 4: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$. (See Figure 8 for an example where the model trained with $N^{\prime}=256$ is highlighted.) Please refer to Figures 10 and 12 for corresponding results with the Llama-3-70B-base model and on the AIME24 dataset, respectively.
@@ 5.4 Approximate DCO also improves MATH answering with chain-of-thought reasoning
Figure 6: Overconfidence persists in CoT fine-tuning with CE loss, which DCOa successfully limits. We fine-tune a Llama-3-8B base model on the MATH dataset on CoT traces. We plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. Left: for CE loss, the model becomes more confident in its most likely answers, as the confidence distribution shifts slightly to the right as training progresses. This effect is milder than in the direct answer setting (Figure 1). Right: for DCOa, the rightward shift in the confidence distribution of the most likely answer over training time is limited; DCOa again limits overconfidence relative to CE loss.
@@ Acknowledgments
We would like to thank Zhinan Cheng, Clémentine Dominé, Marco Fumero, David Klindt, Daniel Kunin, Ben Sorscher and Atsushi Yamamura for helpful discussions. S.G. thanks the James S. McDonnell Foundation, Simons Foundation, NTT Research, Schmidt Foundation and an NSF CAREER Award for support. S.D. and F.C. were partially supported by the McKnight Foundation, the Simons Foundation and an NIH CRCNS grant R01DC020874.
@@ Appendix B Experimental details
All experiments with model size smaller than 10B are performed on machines with 8 NVIDIA H100 GPUs or 8 NVIDIA A100 GPUs. Experiments with model size larger than 10B are performed on machines with 8 NVIDIA H200 GPUs or 8 NVIDIA B200 GPUs. Our codebase uses PyTorch (Ansel et al., 2024), Accelerate (Gugger et al., 2022), and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm (Kwon et al., 2023) for efficient inference. Code will be made available here.
MATH. For experiments with MATH dataset, we fine-tune the Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. In Sections 4 and 5.2 and Figures 1 and 4, we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section 5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
@@
Computational cost. For both DCO and DCOstep, the additional computational cost is minimal—limited to computing the DCO factor, which is negligible compared to the cost of forward and backward passes.
However, DCOa, when applied to fine-tuning with CoT traces, introduces significant overhead due to the need for online Monte Carlo estimation. In the setting of Llama-3 8B, with 32 layers, hidden dimension 4096, context length 1024, and batch size 128, we estimate 5 petaflops for each forward-backward pass. Generating $N_{MC}=64$ CoT’s for each of 128 examples, in turn, requires approximately 50 petaflops. The actual inference compute cost depends on model performance: stronger models, which will discard more samples, will need to process more samples in order to construct a full batch for the next training step. Therefore there is a tradeoff: constructing a batch of 128 hard samples requires running inference on more samples, but also means fewer steps are required to make a full pass through the dataset.
The memory overhead is approximately 30% of GPU memory, when running one inference engine per pair of H100 80GB GPUs. We note that while the flop overhead is significant, costly online inference is also required in RL frameworks for CoT training. Furthermore, the choice of $N_{M}C$ can be made smaller, leading to less precise estimates of the model’s probability of success on a given problem, while speeding up training.
Appendix C Additional results
C.1 Optimality of DCO is achieved when the training objective aligns with the test objective.
Figure 8: Pareto-optimality at a given test strategy pass@N for some $N$ is obtained by DCO training for some $N^{\prime}$ close to $N$. Same as Figure 1 but with $N^{\prime}=256$ highlighted. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.2 Addition results with Llama-3-70B-base
In this section, we present additional results using the larger Llama-3-70B-base model, which are consistent with the findings reported for Llama-3-8B-base.
Table 4: Same as Table 1 but with Llama-3-70B-base model. Pass@N coverage metric on the MATH test set for a Llama-3-70B-base model fine-tuned with CE loss on direct answers from the MATH training set. Surprisingly, Pass@N test accuracy at large $N$ decreases with number of training epochs.
pass@1 pass@16 pass@256 pass@4k
Epoch 1 5.9% 34.3% 68.8% 84.2%
Epoch 2 6.9% 32.0% 63.0% 81.5%
Epoch 3 8.2% 24.2% 48.8% 72.2%
Epoch 4 7.9% 22.2% 44.2% 65.9%
Figure 9: Larger-sized model trained with CE loss also becomes overconfident in its greedy completions; our proposed DCO objective limits this overconfidence on larger-sized models. Same as Figure 1 but with Llama-3-70B-base. We fine-tune a Llama-3-70B-base model on the MATH dataset to produce direct answers without a reasoning trace. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when sampling autoregressively and choosing the most likely token at each step. Leftmost: The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. Left: At the end of the training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up $N$. Right: Same as leftmost but shown for the DCO loss with $N=256$. Relative to the CE loss, the model trained on DCO shows a much milder overconfidence effect. Rightmost: The confidence distribution of the greedy completions after four epochs with DCO for various choices of $N$. As $N$ increases, the model’s confidence on the greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
Figure 10: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier with Llama-3-70B-base. Same as Figure 4 but with Llama-3-70B-base model. We fine-tune Llama-3-70B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.3 Out-of-distribution evaluation
In this section, we present additional evaluation results on AIME24. We confirm that overconfidence persists even on the out-of-distribution test set, and that the DCO algorithm saves test-time scaling by regularizing the max confidence.
Table 5: Same as Table 1 but evaluate on AIME24 dataset. Pass@N coverage metric on the AIME24 dataset for a Llama-3-8B-base model fine-tuned with CE loss on direct answers from the MATH training set. Surprisingly, Pass@N test accuracy at large $N$ decreases with number of training epochs.
pass@1 pass@16 pass@256 pass@4k
Epoch 1 0.2% 3.4% 30.5% 80.5%
Epoch 2 0.3% 4.1% 31.7% 75.6%
Epoch 3 0.2% 2.9% 16.0% 42.0%
Epoch 4 1.9% 4.9% 15.2% 35.8%
Figure 11: A model trained with CE loss becomes overconfident in its greedy completions on out-of-distribution test set; our proposed DCO objective limits this overconfidence on out-of-distribution test set. Same as Figure 1 but evaluated on AIME24. We fine-tune a Llama-3-8B-base model on the MATH dataset to produce direct answers without a reasoning trace and evaluate the model on AIME24. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when sampling autoregressively and choosing the most likely token at each step. Leftmost: The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. Left: At the end of the training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up $N$. Right: Same as leftmost but shown for the DCO loss with $N=256$. Relative to the CE loss, the model trained on DCO shows a much milder overconfidence effect. Rightmost: The confidence distribution of the greedy completions after four epochs with DCO for various choices of $N$. As $N$ increases, the model’s confidence on the greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
Figure 12: DCO improves on CE for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier on out-of-distribution test set. Same as Figure 4 but evaluated on AIME24. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers and evaluate on AIME24. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. We plot pass@N test coverage as a function of $N$, with each curve (solid red or faint blue-green) corresponding to one fine-tuned model. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.4 Theorem proving